OpenStudy (anonymous):

Write y = 2(x - 4)2 + 7 in standard form. Using this standard form of the equation, type the values for a,b and c separated by commas without using spaces in the given box.

OpenStudy (loser66):

open (x-4)^2 first, what do you have?

OpenStudy (anonymous):

x ^2 −8x+16 ?

OpenStudy (loser66):

yup, now combine all, y = 2(x-4)^2 +7 by replace what you have above in.

OpenStudy (loser66):

not yup for your last comment

OpenStudy (anonymous):

i dont understand ?

OpenStudy (anonymous):

a=1, b=−8, c=16

OpenStudy (loser66):

\[y = 2\color{red}{(x-4)^2}+7\\y = 2\color{red}{(x^2-8x+16)}+7\] open the bracket, what do you get?

OpenStudy (anonymous):

(x-4)^2?

OpenStudy (loser66):

a is not 1, you must open the bracket first.

OpenStudy (anonymous):

i need values of a,b,c

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

its not saying anything different

OpenStudy (anonymous):

a=2,b=-8c=16?

OpenStudy (anonymous):

@Loser66 is ryt ,... follow him dude

OpenStudy (loser66):

I know, but!!! you have to follow step by step to get the standard form first, then consider the value of a, b, c You skip the "open the bracket" step, so that, you are wrong.

OpenStudy (anonymous):

i dont know how to do that

OpenStudy (anonymous):

@desirae_michelle17 simply multiply the outer thing to inner things of bracket and add the contants..

OpenStudy (anonymous):

im confused :/ which do i multiply >

OpenStudy (anonymous):

multiply outer 2 with components in bracket

OpenStudy (anonymous):

which bracket ?

OpenStudy (anonymous):

16 & 112?

OpenStudy (anonymous):

y=2(x2-8x+16)+7 y=2x2-16x+32+7 ryt

OpenStudy (anonymous):

now add 32 and 7 that will be easy

OpenStudy (anonymous):

39?

OpenStudy (anonymous):

yup its c 39

OpenStudy (anonymous):

b=16?

OpenStudy (anonymous):

a= ?

OpenStudy (anonymous):

\[2(x-4)^2+7\]=[2(x^2-8x+16)+7\]\[2x^2-16x+32+7\]\[2x^2-16x+39\]

OpenStudy (anonymous):

a=2,b=16,c=39 ?

OpenStudy (anonymous):

b = -16

OpenStudy (anonymous):

okay thank you

OpenStudy (anonymous):

my pleasure

OpenStudy (anonymous):

my pleasure